OpenStudy (o-b-server):

Transforming Equations. Decide whether each equation has one solution, no solutions, or infinitely many solutions.

OpenStudy (o-b-server):

2(x-3) = 2x 3(y-3) = 2y - 9 + y 10x -2 -6x = 3x - 2 + x 4(x+3) + 2x = x - 8